OpenStudy (mtalhahassan2):

For f(x) = 1/2x, determine a) f(1)- f(3) b) f(1/4) + f(3/4)

OpenStudy (anonymous):

Plug in 1, 3, 1/4, and 3/4 into that equation

OpenStudy (anonymous):

When you get 1 and 3. Subtract your answer f(3) from f(1) When you get 1/4 and 3/4, Add you answer for f(1/4) and f(3/4)

OpenStudy (mtalhahassan2):

I get 1/2 - 1/6

OpenStudy (anonymous):

1/2 multiplied by 3 means that you would multiply 3 by 1, but keep the 2 at the bottom. So, A would be 1/2-3/2

OpenStudy (mtalhahassan2):

confuse

OpenStudy (aum):

Is \(\large f(x) = \frac 12x\) or \(\large f(x) = \frac{1}{2x}\) ?

OpenStudy (mtalhahassan2):

the second one is corrcet

OpenStudy (mtalhahassan2):

for the f(1) I get like 1/2

OpenStudy (mtalhahassan2):

and for the f(3) I get 1/6

OpenStudy (aum):

\[ f(x) = \frac {1}{2 x} \\ f(1) = \frac {1}{2*1} = \frac 12 \\ f(3) = \frac {1}{2*3} = \frac 16 \\ f(1) - f(3) = \frac 12 - \frac 16 = \frac 36 - \frac 16 = \frac {3-1}{6} = \frac 26 = \frac 13 \]

OpenStudy (mtalhahassan2):

then I think we have to subtrtact 1/2 - 1/6

OpenStudy (mtalhahassan2):

why did you write 1/3 it be only 3

OpenStudy (aum):

\(\Large \frac 12 - \frac 16\). Find the least common denominator which is 6. Make the denominator 6. So multiply top and bottom of \(\Large \frac 12\) by 3: \(\Large \frac 12 = \frac{1*3}{2*3} = \frac{3}{6} \) \(\Large \frac 12 - \frac 16 = \frac 36 - \frac 16\). If the denominators are the same, the numerators can be combined: \(\Large \frac 36 - \frac 16 = \frac{3-1}{6} = \frac 26 = \frac 13\)

OpenStudy (aum):

Go through some online examples and YouTube videos on adding and subtracting fractions.

OpenStudy (aum):

I have to log out now.

OpenStudy (mtalhahassan2):

ok but what about the part b

OpenStudy (aum):

\[ f(x) = \frac {1}{2 x} \\ f(1/4) = \frac {1}{2*1/4} = \frac {1}{1/2} =2 \\ f(3/4) = \frac {1}{2*3/4} = \frac {1}{3/2} = \frac 23 \\ f(1/4) + f(3/4) = 2 + \frac 23 = \frac{6+2}{3} = \frac 83 \]